### v1

# SteerMusic: Enhanced Musical Consistency for Zero-shot Text-Guided and Personalized Music Editing

Abstract.
Music editing is an important step in music production, which has broad applications, including game development and film production. Most existing zero-shot text-guided methods rely on pretrained diffusion models by involving forward-backward diffusion processes for editing. However, these methods often struggle to maintain the music content consistency. Additionally, text instructions alone usually fail to accurately describe the desired music. In this paper, we propose two music editing methods that enhance the consistency between the original and edited music by leveraging score distillation. The first method, SteerMusic, is a coarse-grained zero-shot editing approach using delta denoising score. The second method, SteerMusic+, enables fine-grained personalized music editing by manipulating a concept token that represents a user-defined musical style. SteerMusic+ allows for the editing of music into any user-defined musical styles that cannot be achieved by the text instructions alone. Experimental results show that our methods outperform existing approaches in preserving both music content consistency and editing fidelity. User studies further validate that our methods achieve superior music editing quality. Audio examples are available on https://steermusic.pages.dev/.
Text-guided music editing, Personalized music editing, Score distillation, Diffusion model
Work done during the internship at Sony AI
copyright: noneccs: Applied computing Sound and music computingccs: Computing methodologies Neural networks
Figure 1. Steering the music style with text-guided music editing or personalized music editing.
1. Introduction
Text-guided diffusion probabilistic models (DPMs) (Ho et al., 2020; Song et al., [n. d.]) have shown impressive performance in generating diverse and high-quality audio samples, including music, speech, and sound effects. Following the success on language and image domain (Wu et al., 2023; Raffel et al., 2020; Rombach et al., 2022), these text-to-audio diffusion models (Liu et al., 2023a, 2024b; Yang et al., 2023; Deepanway et al., 2023; Majumder et al., 2024) trained with huge datasets can generate diverse samples conditioned on the natural language prompts specified. Consequently, the text-guided music editing task (Wang et al., 2023a; Zhang et al., 2024b) was proposed, which performs music editing by modifying the corresponding text prompts of the text-to-audio generation models, offering greater flexibility in applications. Unlike controllable music generation, music editing task modifies an existing piece of music, which revolves around two primary objectives: preserving original musical content and ensuring alignment between the edited music and the desired target.
Previous methods focus on training music editing models from scratch (Agostinelli et al., 2023; Copet et al., 2023; Wang et al., 2023a) or fine-tuning pretrained text-to-audio models (Paissan et al., 2023; Zhang et al., 2024a), both of which require additional datasets or computational costs. Inspired by recent advances in image editing (Brooks et al., 2023; Meng et al., 2021; Huberman-Spiegelglas et al., 2024; Hertz et al., 2022; Zhang et al., 2021), emerging music editing methods have instead pursued zero-shot techniques to reduce computational overhead. Existing zero-shot text-guided music editing approaches (Zhang et al., 2024b; Manor and Michaeli, 2024; Liu et al., 2024a) introduce noise into the source music during the forward diffusion process to suppress high-frequency components (e.g., timbral information) and subsequently perform editing during the denoising phase based on the target guidance. The process of retrieving a noisy latent representation from the data is commonly referred as inversion step. Due to the imperfect diffusion inversion process, the obtained latent representation may fail to fully preserve the original music content. This issue becomes more serious when the source prompt used as the condition cannot accurately capture the detailed characteristics of the music input (Kawar et al., 2023; Paissan et al., 2023). We refer to the distortion in the inverted latent representation as an “inversion error”. Notably, this distortion occurs even in the reconstruction of a few steps of DDIM inversion (Song et al., [n. d.]) , which changes the original melodic information of the reconstructed results as shown in the CQT spectrogram (Brown, 1991) in Figure 2. In music editing, such distortion can even be compounded, resulting in a failure to preserve instruction-irrelevant content in the original music (Zhang et al., 2024a; Kawar et al., 2023; Meng et al., 2021; Brooks et al., 2023). Although methods such as textual inversion (Gal et al., 2022) have been proposed to mitigate this issue by tuning the textual embeddings for near-lossless reconstruction (Niu et al., 2024), there is no way to manipulate target text prompt within the tuned source textual embeddings in editing tasks. A more promising solution to avoid inversion error is delta denoising score (DDS) (Hertz et al., 2023), a score distillation-based method, which performs editing directly in the data space by computing the difference in denoising scores between the source and target prompts through a single forward step. This approach eliminates the dependency of a full or partial forward diffusion process that could introduce inversion errors. By operating in the data space, DDS enables high-fidelity editing while preserving instruction-irrelevant content on the source data.
Figure 2. How inversion error alters the reconstructed melody in CQT spectrogram after 20 steps DDIM inversion.
Text-guided music editing enables flexible and intuitive modifications, requiring users to provide only an arbitrary text instruction to perform the desired edit. However, one limitation for text-guided music editing is that it often lacks fine-grained control over the direction and nuance of editing. For instance, editing a guitar performance into the one played by person A with a guitar brand B. Text-based editing alone struggles to specify the exact “guitar” required. Moreover, person A and guitar brand B might be unseen concepts for the models, rendering attempts to specify these words in the text prompt ineffective. To address this limitation and enhance user-personalized controllability in text-to-music generation, DreamSound (Plitsis et al., 2024) introduces a pioneering approach that adapts image personalization techniques  (Gal et al., 2022; Ruiz et al., 2023) to the music domain, allowing the extraction of user-defined musical characteristics from reference audio. Besides, DreamSound also demonstrates the potential of leveraging the personalization techniques to perform personalized music editing by manipulating learned musical concepts on the noisy source latent through the denoising process of a personalized diffusion model. Despite their attempts, DreamSound still suffers from the inversion error, and struggles to preserve music content while editing the specific concept given in the text prompt.
In this paper, we propose two music editing methods, SteerMusic and SteerMusic+, that can be easily adapted to existing text-to-music DPM based on the score distillation technique. We summarize our key contributions as follows.
  1. (1)
    We propose SteerMusic, a zero-shot text-guided music editing pipeline based on a DDS framework, which focuses on coarse‐level editing, producing high‐fidelity results while preserving source music contents.
  2. (2)
    We propose SteerMusic+, a personalized music editing method that leverages user-defined musical concepts to enable fine-grained customized editing. SteerMusic+, an extension of SteerMusic, enables editing results that are not attainable through text prompts alone. For example, from reggae to the customized reggae given the reference as in Figure 1.
  3. (3)
    We provide extensive experiments to demonstrate that the proposed methods produce superior editing results compared to the existing state-of-the-art methods in terms of musical consistency and edit fidelity.
2. Related work
2.1. Text-guided Music Generation and Editing
In the field of music generation, prior research focuses on approaches that use low-level control signals with strict temporal alignment to the audio outputs. For example, lyrics-conditioned music generation (Yu et al., 2021; Dhariwal et al., 2020) and MIDI-conditioned music generation (Yang et al., 2017). Recently, inspired by advances in natural language processing, high‐level semantic descriptions such as text and visual prompts are gaining popularity in music generation (Dhariwal et al., 2020; Liu et al., 2023a, 2024b; Kundu et al., 2024; Huang et al., 2023a; Chowdhury et al., 2024; Agostinelli et al., 2023; Kreuk et al., 2022; Huang et al., 2023b; Lan et al., 2024; Le Lan et al., [n. d.]; Li et al., 2024; Saito et al., [n. d.]). More recent work also explores melody prompts extracted from reference music (Copet et al., 2024; Novack et al., 2024b, a; Hou et al., 2025; Chen et al., 2024; Kumari et al., 2023), offering more precise and user‐driven music generation capabilities.
Unlike music generation, music editing aims at transforming existing music audio into a target music satisfying the given condition such as text prompts. The goal of music editing is to preserve the original music content while performing the editing. Previous methods involve training music editing models from scratch (Agostinelli et al., 2023; Copet et al., 2023; Wang et al., 2023a; Hou et al., 2025; Mariani et al., 2023) or fine-tuning existing models (Paissan et al., 2023; Zhang et al., 2024a; Tsai et al., 2024; Han et al., 2023; Liu et al., 2023b). Recent methods focus on exploring zero-shot text-guided pipeline (Zhang et al., 2024b; Manor and Michaeli, 2024; Liu et al., 2024a), which rely on pretrained audio generation models. Existing zero-shot text-guided pipelines involve forward-backward diffusion process, which may suffer inversion error as shown in Figure 3 (a).
2.2. Personalized Music Generation and Editing
DreamSound (Plitsis et al., 2024) explores the possibility of capturing musical concepts from the given reference music using a personalization diffusion model (Ruiz et al., 2023; Gal et al., 2022; Kumari et al., 2023). This method allows users to generate new music samples by incorporating the captured personalized musical concept token into the text prompts. In addition to personalized music generation (Plitsis et al., 2024; Chen et al., 2024), DreamSound further extends their method to personalized music editing. However, we noticed that personalized music editing is still immature. Existing methods (Plitsis et al., 2024) struggle to maintain the musical consistency while editing into the desired music concept captured in the reference audio, which use the same music editing pipeline illustrated in Figure 3 (a).
3. Preliminaries
Figure 3. Overview of two branches of music editing pipeline. (a) is existing music editing pipeline, which involves a inversion process conditioned on source prompt, and perform editing on during denoising; (b) is the proposed pipeline where perform editing on data space by optimizing $x=g(\theta)$. [$S$] indicates a concept token defined by users and gray cycles indicate the optimization trajectory from source to target.
3.1. DPMs on Audio Generation
Given an input audio data $x_{0}$ and the corresponding text prompt $y$, the forward diffusion process is defined as a Markov chain that gradually adds noise to the input audio data $x_{0}$ over $T$ steps, and sampled from marginal distribution $q(x_{t}|x_{0}):=\mathcal{N}(x_{t};\sqrt{\alpha_{t}}x_{0},(1-\alpha_{t})I)$ as $x_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}\epsilon_{t}$, where $\alpha_{t}$ represents the variance of the forward process at the time step $t$, $\epsilon_{t}\sim\mathcal{N}(0,I)$ at the time step $t$ and $x_{t}$ denotes the noisy latent representation of the input $x_{0}$ at time step $t$. DPMs (Ho et al., 2020) learn the backward diffusion process, which denoises from the prior distribution $\mathcal{N}(0,I)$ to the data distribution $x_{0}$. By predicting the noise $\epsilon_{\phi}(x_{t},y,t)$, the model parameter $\phi$ can be optimized via the following training objective function:
(1) $\mathcal{L}_{\text{DPM}}(x_{0},y;\phi)=\mathbb{E}_{x_{t}\sim q(x_{t}|x_{0}),t\sim\mathcal{U}(1,T)}\|\epsilon_{\phi}(x_{t},y,t)-\epsilon_{t}\|_{2}^{2}$
where $\epsilon_{t}\sim\mathcal{N}(0,I)$ is the actual noise added to $x_{0}$, $\mathcal{U}(1,T)$ is a uniform distribution. In audio generation, the diffusion process can be either performed in data space or in a latent space obtained by VAEs (Kingma et al., 2013; Rombach et al., 2022). In this study, we perform music editing using AudioLDM2 (Liu et al., 2024b), a latent diffusion model (LDM) pretrained on Mel-spectrograms. During inference, text prompts are given to AudioLDM2 to generate latent samples. The generated samples are decoded by the pretrained VAE decoder back to the Mel-spectrogram. HiFi-GAN (Kong et al., 2020), a commonly used vocoder, converts the Mel-spectrogram into waveforms.
3.2. Score Distillation and Delta Denoising Score
Score distillation is a technique that leverages the score (i.e., the gradient of the log-density) obtained from a pretrained score-based model (e.g., a DPM $\epsilon_{\phi}$) to refine generated samples, which ensures the alignment of predefined constraints. It is typically implemented via probability density distillation, where gradients from the source diffusion model are used to iteratively refine a differentiable function until the desired outcome is achieved. Score distillation sampling (SDS) (Poole et al., 2022) is the pioneering method in score distillation. Given a differentiable function $x=g(\theta)$ that renders a source input data $x$ with parameters $\theta$, SDS optimizes $\theta$ through a differentiable loss $\mathcal{L}_{\text{Diff}}(\phi,x=g(\theta),y^{\text{tgt}})=\mathbb{E}_{t,\epsilon}[w(t)\|\epsilon_{\phi}(x_{t},y^{\text{tgt}},t)-\epsilon\|_{2}^{2}]$. By omitting the UNet Jacobian $\frac{\partial\epsilon_{\phi}}{\partial x_{t}}$ (Poole et al., 2022), the gradient over $\theta$ is given by
(2) $\nabla_{\theta}\mathcal{L}_{\text{SDS}}(\phi,x,y^{\text{tgt}})=\mathbb{E}_{t,\epsilon}[w(t)(\epsilon_{\phi}(x_{t},y^{\text{tgt}},t)-\epsilon)\frac{\partial x}{\partial\theta}]$
where $y^{\text{tgt}}$ is the target prompt, $x_{t}$ is a noisy latent of $x=g(\theta)$ at time step $t$, $w(t)$ is a weighting function. However, SDS is known to suffer from issues such as blurriness in image editing  (Wang et al., 2023b; Hertz et al., 2023). The delta denoising score (DDS) (Hertz et al., 2023) addresses this issue by computing the delta score between the source prompt $y^{\text{src}}$ and the target prompt $y^{\text{tgt}}$. In image editing, DDS refines only the relevant regions of the image corresponding to $y^{\text{tgt}}$, while keeping the other parts of the image unchanged. Given a source input $x^{\text{src}}$ with the corresponding source prompt $y^{\text{src}}$ and the target prompt $y^{\text{tgt}}$, the DDS gradient over the parameters $\theta$ of the differentiable function is
(3) $\nabla_{\theta}\mathcal{L}_{\text{DDS}}(\phi,x,y^{\text{tgt}},x^{\text{src}},y^{\text{src}})=\mathbb{E}_{t,\epsilon}[w(t)(\epsilon_{\phi}(x_{t},y^{\text{tgt}},t)\\ -\epsilon_{\phi}(x^{\text{src}}_{t},y^{\text{src}},t))\frac{\partial x}{\partial\theta}]$
where $x_{t}$ and $x_{t}^{\text{src}}$ shares the same sampled noise $\epsilon$ and time step $t$. See our Supplementary Materials for more detailed background.
While there are different variants of SDS and DDS proposed for image editing (Yu et al., 2025; Hertz et al., 2023; Nam et al., 2024; Lin et al., 2025), the application of DDS in music editing remains underexplored. In the next section, we will explain how to incorporate DDS into music editing.
4. Method
In this section, we introduce two music editing methods. The first method SteerMusic focuses on zero-shot text-guided music editing, and the second method SteerMusic+ focuses on personalized music editing based on a user-defined concept captured from a set of reference music. Unlike the existing forward-backward music editing pipeline in Figure 3 (a), our methods follow a delta denoising framework and perform editing directly on the data space (Figure 3 (b)), which produces results with superior musical content consistency.
4.1. SteerMusic: Zero-shot Text-guided Music Editing
Figure 4. An overview pipeline of SteerMusic+. (a) illustrates a pipeline of fine-tuning a personalized diffusion model given $\mathcal{D}^{\text{ref}}$ and a user-defined [guitar] concept token. (b) provides the detail SteerMusic+ pipeline on personalized music editing that relies on a PDM $\epsilon_{\phi^{\prime}}$ obtained by Eq. 4 in (a). In this example, the PDM captures a user-defined guitar concept. The red dashed lines represent the gradient flows.
We introduce SteerMusic, a zero-shot text-guided music editing method that performs editing in the data space. In this setting, our goal is to edit a source music signal $x^{\text{src}}$ by modifying its corresponding text prompt $y^{\text{src}}$, where $y^{\text{src}}$ is a short textual description of the source music that includes the specific musical attribute intended for modification. The modified text prompt $y^{\text{tgt}}$ acts as the target prompt to guide the editing of $x$. To obtain desirable edited results, the musical content shared by both $y^{\text{tgt}}$ and $y^{\text{src}}$ should be preserved, changing only the contents that are distinct in the $y^{\text{tgt}}$. For example, if the only change in $y^{\text{tgt}}$ compared to $y^{\text{src}}$ is replacing only the word “piano” with “guitar” while keeping the rest of the sentence unchanged, the edited music should preserve the melody and tempo, while modifying only the instrument sound.
To achieve this goal, we adopt the DDS method (Hertz et al., 2023), which has previously been explored only in the image domain. To the best of our knowledge, we are the first to investigate the potential application in music editing. Given a differentiable mapping function $x=g(\theta)$, where we initialize $x=x_{0}^{\text{src}}$, the gradient over $\theta$ is the same as in Eq. 3. In SteerMusic, we set the pretrained diffusion model $\epsilon_{\phi}$ as AudioLDM2 (Liu et al., 2024b). Similar to (Hertz et al., 2023), the delta score in Eq. 3 steers the optimization process toward the target prompt while reducing the noisy editing direction commonly associated with vanilla SDS, leading to enhanced edit fidelity. SteerMusic method enables flexible editing using text instructions alone; however, it is limited to coarse-grained editing, as text prompts often lack the precision information to capture fine-grained musical details.
4.2. SteerMusic+: Personalized Music Editing
As mentioned in Sec. 2, text-guided editing lacks customization for precise music editing, such as transferring music to a specific style. To enable fine-grained personalized editing, we propose SteerMusic+, an extension of SteerMusic. In this setting, we have a set of source music and prompt pair $\{x^{\text{src}},y^{\text{src}}\}$. The target prompt $y^{\text{tgt}}$ is constructed by modifying $y^{\text{src}}$ to manipulate a user-defined concept token $[S]$, representing the desired customization (see example in Figure 4 (b)). SteerMusic+ uses two pretrained diffusion models:
  • A pretrained diffusion probabilistic model (DPM), denoted as $\epsilon_{\phi}$, which serves as a reference for maintaining consistency with the original music content; and
  • A personalized diffusion model (PDM), denoted as $\epsilon_{\phi^{\prime}}$, which is fine-tuned on a small set of dataset contains reference music to capture the user-defined style concept $[S]$ and guide the editing process toward the desired direction.
Inspired by (Yu et al., 2025), we define successful personalized editing based on the following criteria:
  • The instruction-irrelevant part (i.e., {$y^{\text{tgt}}\cap y^{\text{src}}$}) in the source music should be maintained.
  • The edited musical attributes should perceptually align with the intended personalized musical concept $[S]$.
In the following parts, we present SteerMusic+ method that enables personalized music editing with enhanced music consistency.
Personalized Diffusion Model (PDM) serves a foundational part in SteerMusic+. Since training a PDM has been extensively studied in (Plitsis et al., 2024; Ruiz et al., 2023; Gal et al., 2022; Kumari et al., 2023), we assume the availability of a pretrained text-to-music PDM in SteerMusic+, as SteerMusic+ is a plug-in pipeline compatible with existing pretrained PDM. The text-to-music PDM $\epsilon_{\phi^{\prime}}$ captures the user-defined musical concept $S$ by fine-tuning an existing pretrained DPM $\epsilon_{\phi}$ under a small set of reference music $\mathcal{D}^{\text{ref}}=\{(x^{\text{ref}},y^{\text{ref}})_{n}\}^{N}_{n=1}$, where $N$ can be as few as $1$ (Plitsis et al., 2024). The fine-tuning is achieved via optimizing the objective
(4) $\phi^{\prime}\in\text{arg min}_{\hat{\phi}}\mathbb{E}_{(x^{\text{ref}},y^{\text{ref}})\sim\mathcal{D}^{\text{ref}}}\mathcal{L}_{\text{DPM}}(x^{\text{ref}},y^{\text{ref}};\hat{\phi})$
where $\hat{\phi}$ is initialized with a pretrained DPM weights $\phi$. As illustrated in Figure 4 (a), the prompt $y^{\text{ref}}$ takes form of “a recording of a [$S$]”, where the placeholder [$S$] corresponds to a defined new concept word embedding. During inference, the PDM can generate music with the newly learned concept (e.g., “A disco song with a [$S$]”). The dataset $\mathcal{D}^{\text{ref}}$ consists of reference audios that encapsulate the concept users aim to extract. According to (Plitsis et al., 2024), the concept represents a musical style, which can be either an instance of instrument sounds or a specific genre that cannot be yielded even with the most detailed textual description. For instance, if the objective is to capture the user’s guitar playing style, the reference audio should feature performances on the user’s guitar playing. Conversely, if the goal is to capture the concept of jazz, the reference audio should consist of recordings that exemplify the same jazz style.
Although PDM can generate new music compositions with new concepts captured from the reference data set $\mathcal{D}^{\text{ref}}$, it cannot modify existing music using $y^{\text{tgt}}$ containing the new concepts directly. Next, we introduce the key components of SteerMusic+ that enable the PDM to perform personalized music editing.
Personalized Delta Score (PDS) is an extension of Eq. 3 to enable personalized music editing. We define the PDS loss as $\mathcal{L}_{\text{PDS}}(\phi^{\prime},\phi,\\ x,y^{\text{tgt}},x^{\text{src}},y^{\text{src}})=\mathbb{E}_{t,\epsilon}[w(t)\|\epsilon_{\phi^{\prime}}(x_{t},y^{\text{tgt}},t)-\epsilon_{\phi}(x^{\text{src}}_{t},y^{\text{src}},t)\|^{2}_{2}]$. By omitting the UNet Jacobian, the gradient over $\theta$ is given by
(5) $\nabla_{\theta}\mathcal{L}_{\text{PDS}}=\mathbb{E}_{t,\epsilon}[w(t)(\epsilon_{\phi^{\prime}}(x_{t},y^{\text{tgt}},t)-\epsilon_{\phi}(x^{\text{src}}_{t},y^{\text{src}},t))\frac{\partial x}{\partial\theta}]$
where $x_{t}$ and $x_{t}^{\text{src}}$ shares the same sampled noise $\epsilon$ at time step $t$.
This modified delta score can be decomposed to two components: The score of $\epsilon_{\phi^{\prime}}$ provides a desired direction to guide the editing to match the target prompt with concept $[S]$. The score of $\epsilon_{\phi}$ helps to reduce noisy direction of unintended modification areas. The delta score between the PDM $\epsilon_{\phi^{\prime}}$ and the DPM $\epsilon_{\phi}$ may not yield an effective direction toward $y^{\text{tgt}}$, as $\epsilon_{\phi^{\prime}}$ shifted to the reference distribution $\mathcal{D}^{\text{ref}}$. We introduce an additional component to compensate for the distribution shift induced by the score of $\epsilon_{\phi^{\prime}}$.
Distribution Shift Regularization. To bridge the distribution gap between $\epsilon_{\phi^{\prime}}$ and $\epsilon_{\phi}$, we introduce a regularization term to further regularize the edited score to the personalized diffusion model $\epsilon_{\phi^{\prime}}$. We wish to minimize the distribution shift between two diffusion models by adding a constraint as
(6) $\min_{\theta}$ $\mathcal{L}_{\text{PDS}}(\phi^{\prime},\phi,x,y^{\text{tgt}},x^{\text{src}},y^{\text{src}}),$
subject to $\mathcal{L}_{\text{shift}}(\phi^{\prime},\phi,x,y^{\text{tgt}})-\zeta\leq 0.$
where $\zeta$ is a small amount of constant, the regularization term is $\mathcal{L}_{\text{shift}}(\phi^{\prime},\phi,x,y^{\text{tgt}})=\mathbb{E}_{t,\epsilon}[w(t)\|\epsilon_{\phi^{\prime}}(x_{t},y^{\text{tgt}},t)-\epsilon_{\phi}(x_{t},y^{\text{tgt}},t)\|^{2}_{2}]$. Same as above, the gradient with respect to $\theta$ is given by
(7) $\nabla_{\theta}\mathcal{L}_{\text{shift}}=\mathbb{E}_{t,\epsilon}[w(t)(\epsilon_{\phi^{\prime}}(x_{t},y^{\text{tgt}},t)-\epsilon_{\phi}(x_{t},y^{\text{tgt}},t))\frac{\partial x}{\partial\theta}]$
By integrating the constraint, the overall gradient through $\theta$ is
(8) ${}\nabla_{\theta}\mathcal{L}_{\text{PDS-O}}(\phi^{\prime},\phi,x,y,x^{\text{src}},y^{\text{src}})=\underbrace{\nabla_{\theta}\mathcal{L}_{\text{PDS}}(\phi^{\prime},\phi,x,y^{\text{tgt}},x^{\text{src}},y^{\text{src}})}_{\text{`Delta score point to edit direction'}}\\ +\lambda\underbrace{\nabla_{\theta}\mathcal{L}_{\text{shift}}(\phi^{\prime},\phi,x,y^{\text{tgt}})}_{\text{`Delta score regularizes distribution shift'}}$
where $\lambda$ is a predefined constant that adjusts the strength of the regularization. We will discuss the effect of $\lambda$ in our ablation study.
Table 1. Model comparison on zero-shot text-guided music editing task.
[…]
DDIM (Song et al., [n. d.]) 0.477 1.022 0.330 5.377 0.264 1.37 1.91
SDEdit (Meng et al., 2021) 0.638 1.274 0.169 6.208 0.218 0.92 1.68
MusicMagus (Zhang et al., 2024b) 0.593 1.698 0.338 5.243 0.238 2.11 1.57
ZETA (Manor and Michaeli, 2024) 0.509 1.021 0.293 5.458 0.252 1.22 1.60
SteerMusic (Ours) 0.278 0.381 0.480 3.772 0.259 2.92 2.50
Personalized Contrastive (PCon) Loss. Eq. 8 formulates a regularized editing direction, where the regularized delta score encourages alignment with the target prompt while mitigating the distribution shift. It does not explicitly enforce fidelity to the concept, which may lead to suboptimal editing quality. To further enhance the edit fidelity, we incorporate a PCon loss between temporal features, which is modified from a patch-wise contrastive loss (Nam et al., 2024). PCon loss extracts intermediate features $h^{\text{src}}_{l}$ and $h_{l}$ that passed through the residual block and self-attention block from $\epsilon_{\phi}$ conditioned on $y^{\text{src}}$ and $\epsilon_{\phi^{\prime}}$ conditioned on $y^{\text{tgt}}$ at $l$-th self-attention layer, respectively. The features are then reshaped into size $\mathbb{R}^{T_{l}\times F_{l}\times C_{l}}$, where $T_{l},F_{l}$, and $C_{l}$ represent the size of the temporal, spatial, and channel dimension in the $l$-th layer, respectively. The patch corresponding to the temporal location on the feature map $h^{\text{src}}_{l}$ is designated as ‘positive’, and vice versa. The PCon loss is defined as
(9) $\mathcal{L}_{\text{PCon}}(x,x^{\text{src}})=\mathbb{E}_{h}[\sum_{l}\sum_{t^{\prime}}\ell(h_{l}^{t^{\prime}},h_{l}^{\text{src},t^{\prime}},h_{l}^{\text{src},T_{l}\backslash t^{\prime}})]$
(10) $\ell(h,h^{+},h^{-})=-\text{log}(\frac{\text{exp}(h\cdot h^{+}/\tau)}{\text{exp}(h\cdot h^{+}/\tau)+\text{exp}(h\cdot h^{-}/\tau)})$
where $t^{\prime}\in\{1,...,T_{l}\}$ represents temporal location query patch, the positive patch as $h_{l}^{src,t^{\prime}}$ while the other patches as $h_{l}^{src,T_{l}\backslash t^{\prime}}$. $\text{exp}(h\cdot h^{+}/\tau)$ is positive sample with the same temporal location, $\text{exp}(h\cdot h^{-}/\tau)$ is negative sample with mismatched temporal location in the self-attention features, $\tau$ is a temperature parameter as $\tau>0$.
The gradient of $\mathcal{L}_{\text{PCon}}(x,x^{\text{src}})$ will propagate to the hidden state of self-attention layers $h$ in the personalized diffusion $\epsilon_{\phi^{\prime}}$. Given that the personalized diffusion $\epsilon_{\phi^{\prime}}$ has a distribution shift over the reference dataset $\mathcal{D}^{\text{ref}}$, the $\mathcal{L}_{\text{PCon}}$ explicitly encourages feature similarity at the frequency domain in self-attentions, particularly for attributes that distinguish the target concept. This reinforcement leads the model to prioritize concept consistency over strict temporal alignment with the source music. Consequently, $\mathcal{L}_{\text{PCon}}$ amplifies the distinctive characteristics of the target concept of $\epsilon_{\phi^{\prime}}$ in SteerMusic+, ensuring that the edited music maintains stronger fidelity to the desired style while allowing structural variations.
5. Experiment
5.1. Experimental Setup
To evaluate our methods, we use a pretrained AudioLDM2 (Liu et al., 2024b) as the foundational model. In the text-guided music editing experiment, we set the optimization iteration as 400 with a 30 guidance scale (Ho and Salimans, [n. d.]) in the diffusion model on Eq. 3 in SteerMusic. In personalized music editing experiment, we follow the setting of the official repository of DreamSound [^1] https://github.com/zelaki/DreamSound and fine-tune the DreamSound model 100 steps with $1e^{-5}$ learning rate. We set 400 optimization steps with 15 guidance scale and 0.05 $\lambda$ value in Eq. 8 of SteerMusic+. All experiments were performed on a single NVIDIA H100 GPU.
Figure 5. A top CQT feature comparison between source and text-guided edited music for different methods. SteerMusic (ours) produces edited results have similar top CQT compare to other methods, indicating SteerMusic successfully preserves source music melody.
Figure 6. Radar chart of model comparison for each text-guided music editing subtask on SteerMusic.
5.2. Evaluation Metrics
We evaluate music editing objectively based on two aspects: musical consistency (content preservation before and after editing) and editing fidelity. We follow (Manor and Michaeli, 2024) and calculate the following objective metrics for measurement. To evaluate musical consistency, we use:
Fréchet Audio Distance (FAD) (Kilgour et al., 2019) measures the distributional difference between source and edited music (lower the better). We calculate FAD based on both VGGish (Hershey et al., 2017) and clap-laion-music (Wu et al., 2023) embeddings. LPAPS (Iashin and Rahtu, 2021), an audio version of LPIPS (Zhang et al., 2018) which quantifies the consistency of the edited audio relative to the source audio (lower the better). Top-1 Constant-Q Transform Pearson Correlation Coefficient (CQT-1 PCC) measures melody consistency between the source and edited music (higher the better). The CQT-1 (Brown, 1991) extracts main melody of the music audio and has been shown to outperform traditional chroma-based features in representing melodic characteristics (Hou et al., 2025). While existing metrics such as FAD and LPAPS provide insight into audio quality and perceptual similarity, they fall short in comprehensively capturing melodic structure. Although existing deep learning models (Cwitkowitz et al., 2024; Bittner et al., 2022; Gardner et al., 2022; Chang et al., 2024; Mancusi et al., 2025) effectively extract melodies from real music, they are unreliable for synthesized audio. To address the limitation, we introduce CQT-1 PCC as a supplementary objective metric, specifically designed to quantify melodic consistency in generative music editing. This metric enables a more targeted evaluation of whether the core melodic structure of the source audio is retained in the edited result. More details of the metrics are provided in the Supplementary Material.
To evaluate editing fidelity, we use:
CLAP Score (Wu et al., 2023; Chen et al., 2022) measures alignment between edited music and the target prompt in text-guided music editing (higher is better). CDPAM (Manocha et al., 2021), a perceptual audio metric leveraging deep learning representations, measures perceptual distance between audios such as music and speech (Jacobellis et al., 2024; Hai et al., 2024; Niu et al., 2024; Gui et al., 2024; Chung et al., 2025). We use CDPAM to evaluate audio perceptual similarity between reference music and the edited result in personalized music editing (lower the better).
Subjective evaluation: To further validate our methods, we designed a Mean Opinion Score (MOS) study to evaluate both target editing fidelity (MOS-T) and source content preservation (MOS-P) subjectively by asking participants to rate results from 1-Bad to 5-Excellent (ITU-T RECOMMENDATION, 1999; Sector, 1996) for randomly selected edited samples.222 This study has been reviewed by the internal privacy committee to ensure compliance with local regulations for ethical conduct of research with human participants and ACM’s Code of Ethic. More details are provided in the Supplementary Material.
5.3. Zero-shot Text-guided Music Editing
In this part, we evaluate our SteerMusic method proposed in Sec. 4.1 on zero-shot text-guided music editing task.
Dataset: We use the ZoME-Bench (Liu et al., 2024a) dataset [^3] https://huggingface.co/datasets/liuhuadai/ZoME-Bench to evaluate our method, which contains 1000 10-second audio samples obtained from MusicCaps (Agostinelli et al., 2023), with corresponding source and target prompts for the editing. ZoME-Bench has 9 subdatasets that target different music editing tasks. Since SteerMusic primarily focuses on enhancing musical consistency for music style editing, we select specific subsets from the dataset that involve instrument changes, genre changes, mood changes, and background modifications. These subsets contain 131, 134, 100, and 95 music clips, respectively.
Baseline: We directly compare SteerMusic with other plug-in zero-shot text-guided music editing methods, SDEdit (Wang et al., 2023a; Meng et al., 2021), DDIM (Song et al., [n. d.]), ZETA (Manor and Michaeli, 2024) and MusicMagus (Zhang et al., 2024b), which based on the same pretrained AudioLDM2 (Liu et al., 2024b). We do not compare with MEDIC (Liu et al., 2024a) and MelodyFlow (Le Lan et al., [n. d.]) as their source code has not been released. We also exclude AudioEditor (Jia et al., 2025) and AudioMorphX (Liang et al., 2024), which are primarily designed for sound editing rather than music editing.
Table 2. Model comparison on personalized music editing task.
[…]
Textual inv. (Plitsis et al., 2024) 0.789 1.871 0.216 5.083 0.713 1.64 1.63
DreamSound (Plitsis et al., 2024) 0.902 3.175 0.292 5.082 0.609 1.42 1.81
SteerMusic+ (Ours) 0.362 0.558 0.399 4.125 0.593 3.07 2.47
Figure 7. User preference study for SteerMusic. We report the percentage of users that preferred SteerMusic (ours) over ZETA (Manor and Michaeli, 2024) and MusicMaugs (Zhang et al., 2024b) in this user study.
Figure 8. A visualization of edited results between SteerMusic+ and baselines in personalized music editing. SteerMusic+ preserves instruction-irrelevant musical content on the source music.
Experimental results: Table 1 compares SteerMusic with other zero-shot text-guided music editing baselines across various music style transfer tasks. Our method demonstrates notable improvements in musical consistency, as evidenced by higher CQT-1 PCC, lower LPAPS, and lower FAD scores. DDIM achieves the best CLAP score (0.264), it has a much lower CQT-1 PCC and a significantly higher LPAPS. This suggests that while DDIM achieves higher alignment with the target style in terms of text-audio correspondence, it directly conditions on the target prompt during the denoising process without applying constraints to preserve the original music information. In contrast, SteerMusic balances both source music consistency and target style fidelity, ensuring that the resulting music remains faithful to the original music while effectively incorporating the target style changes. Moreover, subjective evaluation scores (MOS-P and MOS-T) further support the superiority of SteerMusic in both content preservation and edit fidelity, aligning well with the objective metrics. Figure 5 presents a visualized comparison of top-1 CQT features for an instrument editing sample. SteerMusic generates results with CQT features that closely resemble those of the original music, suggesting that it effectively preserves the original music melody consistency. A more detailed comparison across individual subtasks is in Figure 6.444See our Supplementary Material for extra comparison with SDS (Poole et al., 2022) and CDS (Nam et al., 2024).
User preference study: We evaluate SteerMusic by a user preference study following the design in (Manor and Michaeli, 2024)2. To reduce cognitive load and improve reliability, we compare SteerMusic with the top-performing baselines, MusicMagus and ZETA. In this study, users were asked to answer a sequence of 20 questions, each question contains original music, an editing instruction, and two edited results. Users were instructed to select the edited result that matches the instruction better while preserving the main original music content. We collected 25 full responses, which the participants having a minimum of 1 and average of 5 years of music training. Our method was clearly preferred over all competing methods (see Figure 7).
5.4. Personalized Music Editing
In this part, we evaluate our SteerMusic+ method proposed in Sec. 4.2 on personalized music editing task.
Dataset: We selected a representative subset of eight musical concepts from the 32 concepts provided by (Plitsis et al., 2024), for both common and uncommon cases, which include four instrument concepts (i.e., Guitar, Bouzouki, Ocarina, Sitar) and four genre concepts (i.e., Morricone, Reggae, Hiphop, Sarabande). Each musical concept includes an instruction of concept placeholder and five 10-second reference audio clips sourced from YouTube and FreeSound. For the source music on the personalized editing, we follow Sec. 5.3 and use the “change instruments” and “change genre” subsets in ZoME-banch (Liu et al., 2024a) to perform personalized instrument transfer and personalized genre transfer, which contains 131 and 134 audio clips respectively. We manipulate the concept token by adapting the original target prompt provided by ZoME-banch for each audio sample from the original word replacement to the musical concept.
Baseline: We set two existing personalized music editing methods proposed by (Plitsis et al., 2024) as the baselines, Textual inversion and DreamSound. Textual inversion optimizes a concept embedding, whereas DreamSound fine-tunes a AudioLDM2 with rare-token identifiers (Ruiz et al., 2023). Both methods perform personalized music editing by manipulating the concept token during denoising process. We follow the official codes provided by (Plitsis et al., 2024) to obtain text-to-music PDM1. According to (Plitsis et al., 2024), we reproduce the personalized music editing methods by calculating noisy latent representations $x_{t}$ from $x_{0}^{\text{src}}$ of a DPM conditioned on the source prompt with a predefined shallow time step $t$ using DDIM inversion (Song et al., [n. d.]), where $t=30$. We denoise $x_{t}$ on the PDM [^5] Since textual inversion optimizes only the concept token embedding rather than fine-tuning a PDM, we denoise $x_{t}$ using the DPM employed during DDIM inversion. Consequently, textual inversion is not compatible with SteerMusic+. conditioned on the target prompt linked to the learned concept to obtain $x_{0}^{\text{tgt}}$. We do not compare with Jen-1 DreamStyler (Chen et al., 2024), as it focuses on personalized music generation rather than editing.
Experimental results: We plug-in SteerMusic+ in the same PDM used in DreamSound5. Table 2 shows that SteerMusic+ achieves superior musical consistency compare to the two baselines. It indicates that the edited outputs of SteerMusic+ successfully preserve instruction-irrelevant music content in the source music. Furthermore, SteerMusic+ performs accurate edits that align well with the concepts captured from references as indicated by the low CDPAM compared to the baseline methods. These objective results are aligned to the subjective metrics, where SteerMusic+ obtains higher MOS-P and MOS-T than the baselines. Figure 8 provides a visual comparison on personalized music instrument transfer task. SteerMusic+ successfully preserves instruction-irrelevant music content on the source music, such as vocal information. More visualizations are available in our supplementary material.
Figure 9. User preference study for SteerMusic+. We report the percentage of users that preferred SteerMusic+ (ours) over DreamSound and Textual inv. (Plitsis et al., 2024) in this user study.
User preference study: Similar to Sec. 5.3, we evaluate SteerMusic+ by a user preference study compared to DreamSound and Textual inversion2. In this study, users were asked to select the edit result that the style better matches the reference music style while preserving the main source music content. We collect 24 full responses, which participants have minimum of 1 and on average 5 years of music training experience. As shown in Figure 9, SteerMusic+ are clearly preferred by users than the baselines.
Figure 10. Ablation study on SteerMusic+ with adherence to music consistency vs. edit fidelity on the edited music with different $\lambda$ values in Eq. 8. The horizontal axis (CQT-1 PCC) measures how well the source melody is preserved, while the vertical axis (CDPAM) reflects closeness to the target concept.
Ablation study: We conduct an ablation study to understand the effects of different components in SteerMusic+. We used the concept [bouzouki] as it is an uncommon instrument style that typically requires personalized models. Yellow dots indicate results obtained by $\mathcal{L}_{\text{PDS-O}}$ in Eq. 8 alone, with varying $\lambda$ values. From the zoomed-in view in Figure 10, we observe that as the CDPAM value decreases with $\lambda$, the loss of $\mathcal{L}_{\text{shift}}$ within $\mathcal{L}_{PDS-O}$ helps steer the editing process toward the personalized concept. However, the effect of $\mathcal{L}_{\text{shift}}$ is minor, suggesting that the gradient of $\mathcal{L}_{\text{PDS}}$ alone does not sufficiently capture the target concept during editing. As a result, the edited music struggles to align with the reference musical characteristics. In contrast, SteerMusic+ (with $\mathcal{L}_{\text{PCon}}$ in Eq. 9) leads to a significant decrease in CDPAM value, indicating better alignment with the target concept. Since $\mathcal{L}_{\text{PCon}}$ was proposed to explicitly enhance the characteristics of the target concept on the editing, we observe that $\mathcal{L}_{\text{shift}}$ becomes more effective in this setting. Specifically, when $\lambda$ is negative, the edit preserves the original musical content, while positive $\lambda$ push the edit toward the target concept.
This reveals a fundamental trade-off between music consistency and adherence to the target style. In personalized music editing, where musical characteristics are derived from reference music, shifting the edit closer to the concept inevitably disrupts the original content. This is distinct from text-guided music editing, where the transformation follows a textual description rather than an explicit reference. In personalized editing, the challenge lies in balancing the retention of the original musical structure, while integrating the distinctive attributes of the reference music. We suggest that the $\lambda$ value should not be too large for a further regularization in real applications (e.g., $\lambda=0.05$).
Figure 11. Adherence to music consistency vs. edit fidelity to edited results with different fine-tune steps in PDM.
Limitation. As in shown Figure 11, we observe that the number of fine-tune steps with a fixed learning rate in the PDM significantly influences the ability of DreamSound and SteerMusic+ to capture the concept of reference music while preserving the original melody. When the PDM uses few fine-tuning steps (e.g., 50), it struggles to fully internalize the reference concept, resulting in higher CDPAM scores and underwhelming edit consistency. Conversely, with an excessively large number of steps (e.g., 200), the PDM tends to overfit, leading to lower CQT-1 PCC and a loss of source music structure. These findings highlight the importance of balancing fine-tuning steps in the PDM to achieve both high edit fidelity and source content preservation on personalized music editing setting.
6. Conclusion
We present two music editing methods, SteerMusic and SteerMusic+, from coarse-grained to fine-grained music editing. To the best of our knowledge, this is the first attempt to fully explore DDS and score distillation technique in the music editing framework. Our methods overcome the limitation of existing methods that lack musical consistency, causing the edited results to differ from the original music composition. In additional to better musical consistency, our methods also produce high-fidelity edited results that align with the target text prompts. We also extended our method beyond text-guided method and proposed a novel pipeline for personalized music editing. This method allows for fine-grained music editing, which extracts the user-defined music style concept token from reference music, producing more precise and accurate editing. We systematically evaluate our method through quantitative and qualitative metrics, complemented by comprehensive ablation studies. In our experiment, we demonstrated that our methods outperform existing baseline methods, which improves the musical consistency while maintaining high editing fidelity.